a utility pole breaks and falls. nine feet of the pole remains standing ant the top of the pole lands 12 feet from the base forming a right triangle. find the original height of the utility pole.
Answer by checkley77(12844)

9^2+12^2=X^2
81+144=X^2
X^2=225
X=SQRT225
X=15 LENGTH OF THE POLE THAT FELL.
9+15=24 FT. FOR THE POLE HEIGHT.